OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 12 and a13 = −60? @ganeshie8 −192 −186 −180 −174

ganeshie8 (ganeshie8):

whats the general term of arithmetic sequence ?

ganeshie8 (ganeshie8):

\(a_n = a_1 + (n-1)d\)

OpenStudy (anonymous):

12+(32-1)d

ganeshie8 (ganeshie8):

yes, we need d :) we're given :- \(a_1 = 12\) \(a_{13} = -60\) \(-60 = 12 + (13-1)d\) find \(d\) first

ganeshie8 (ganeshie8):

^^solve d

OpenStudy (anonymous):

124d=-60 am i in the right direction

ganeshie8 (ganeshie8):

how ?

ganeshie8 (ganeshie8):

\(-60 = 12 + (13-1)d\) \(-60 = 12 + (12)d\) \(-72 = (12)d\) solve d

OpenStudy (anonymous):

d(13-1) i used distributive property then minused and they equaled 12d

OpenStudy (anonymous):

-6? @ganeshie8

ganeshie8 (ganeshie8):

yes ! d = -6

OpenStudy (anonymous):

ok now i just do that for 32

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

so, we have \(a_1 = 12\) and \(d = -6\) we can find \(a_{32}\) using the general term formula ?

OpenStudy (anonymous):

whats the formula

ganeshie8 (ganeshie8):

\(a_n = a_1 + (n-1)d\) \(a_{32} = a_1 + (32-1)d\)

OpenStudy (anonymous):

a1 is 12 and -6 is d

ganeshie8 (ganeshie8):

plug them and simplify

OpenStudy (anonymous):

i got -2.583 i know that is wrong :(

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(a_n = a_1 + (n-1)d\) \(a_{32} = a_1 + (32-1)d\) \(a_{32} = a_1 + (31)d\)

ganeshie8 (ganeshie8):

plugin a1 and d values, and show me next step :)

ganeshie8 (ganeshie8):

step by step ok

OpenStudy (anonymous):

a_32=12+(31d)

OpenStudy (anonymous):

then i minused 12 from both sides

ganeshie8 (ganeshie8):

nope. first plugin the value for d also

OpenStudy (anonymous):

12-186

OpenStudy (anonymous):

a_32=-174

OpenStudy (anonymous):

!!!omg I think I got it

ganeshie8 (ganeshie8):

\(a_{32} = a_1 + (31)d\) plugin a1 and d values above \(a_{32} = 12 + (31)(-6)\) \(= 12 - 186\) \(=-174\)

ganeshie8 (ganeshie8):

Yes :)

OpenStudy (anonymous):

You are awesome thank you!!!11

ganeshie8 (ganeshie8):

np you're awesome too :D